OpenStudy (anonymous):

On a certain project the only grades awarded were 75 and 100. If 85 students completed the project and the average of their grades was 85, how many earned 100?

OpenStudy (anonymous):

tips, yummy?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

hello patrick

OpenStudy (anonymous):

x=# of students who got 75 y=# of students who got 100 x+y=85 that much we know

OpenStudy (anonymous):

x+y=85

OpenStudy (anonymous):

and 75x+100y=85

OpenStudy (anonymous):

so just solve for y :)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

x+y=85 75x+100y=85 solve for y

OpenStudy (anonymous):

100y=85-75x

OpenStudy (anonymous):

y=17/20-3/4x

OpenStudy (anonymous):

y=7

OpenStudy (anonymous):

x=78?

OpenStudy (anonymous):

so seven people go hundreds?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

x=78

OpenStudy (anonymous):

y=7

OpenStudy (anonymous):

those values are correct, right?

OpenStudy (anonymous):

nope :\ that would make the average 77 or 78, not 85

OpenStudy (anonymous):

im kinda confused myself now

OpenStudy (anonymous):

yummu

OpenStudy (anonymous):

lol wat?

OpenStudy (anonymous):

t carrol is coming

OpenStudy (anonymous):

@jim_thompson5910 or @Hero

OpenStudy (anonymous):

tips carrol?

OpenStudy (anonymous):

85 = N(75) + (85-N)(100). So, 85-N got a 100.

OpenStudy (anonymous):

Just solve for N and hence, 85-N.

OpenStudy (anonymous):

N=336.6?

OpenStudy (anonymous):

wah?

jimthompson5910 (jim_thompson5910):

yummydum's first equation is correct

jimthompson5910 (jim_thompson5910):

the second equation is missing something

jimthompson5910 (jim_thompson5910):

75x+100y represents the sum of all the scores, you need to divide that sum by 85 since there are 85 scores

OpenStudy (anonymous):

x+y=85

jimthompson5910 (jim_thompson5910):

so the second equation should be (75x+100y)/85 = 85

OpenStudy (anonymous):

(75x+100y)/85=85

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

solve for y?

jimthompson5910 (jim_thompson5910):

yes, y = # of people who got 100

OpenStudy (anonymous):

y==71.5

OpenStudy (anonymous):

y=71.4

OpenStudy (anonymous):

y=71.5

jimthompson5910 (jim_thompson5910):

no, you should get a whole number

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

75x+100y=85 squared?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

y=72.25-3xy?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

10 square root of 75=85?

jimthompson5910 (jim_thompson5910):

x+y=85 x=85-y ----------------- 75x+100y=85^2 75x+100y=7225 75(85-y)+100y=7225 solve for y

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

I did it too fast the first time, sorry. I re-did it and got it right. If you don't get the answer from the other helpers, let me know and I can show you.

OpenStudy (anonymous):

y=34

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

that's the answer

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yes. that's what I also got.

OpenStudy (anonymous):

these problems are from the PSAT

OpenStudy (anonymous):

I want to solve them efficiently in order to do well in the test

OpenStudy (anonymous):

I want to get at least a 200 on my PSAT

OpenStudy (anonymous):

You're on your way. Good luck on the PSAT.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

see you guys later